The food compartment of a refrigerator, is maintained at 4°C by

removing heat from it at a rate of 360 kJ/min. If the required power input

to the refrigerator is 2 kW, determine (a) the coefficient of performance of

the refrigerator and (b) the rate of heat rejection to the room that houses

the refrigerator.


1
Expert's answer
2021-01-19T02:32:12-0500

(a) the coefficient of performance of the refrigerator


"{\rm COP}=\frac{P_C}{P}=\rm\frac{360\times \frac{1000\: J}{60\: s}}{2000\: W}=3"

(b) the rate of heat rejection to the room that houses the refrigerator

"P_H=P_C+P=\rm 360\times \frac{1000\: J}{60\: s}+2000\: W=8000\; W"